A transparent plastic has been developed. When light travels through it, it only travels at 3 × 107 m/s. What will be the refractive index of this new plastic?


1
Expert's answer
2021-07-19T09:51:48-0400

"n=\frac cv,"

"n=\frac{3\cdot 10^8}{3\cdot 10^7}=10."